If the measure of two complementary angles are in the ratio 4:1 what is the smaller ratio?
Answer by rapaljer(4551)

Let x=smaller angle
4x= larger angle

x+4x=90
5x=90
x=18 degrees
4x=4*18=72 degrees

Check: 18+72=90

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus